3.15 Project Management What Is the Project Duration in Weeks

Document Sample
3.15 Project Management What Is the Project Duration in Weeks Powered By Docstoc
					IT Services Weekly Project Status Report
6-Mar-06



                                                                                                                                                                                                                                                                                   Hot Issues
                                                                           Initia Project        Project                                                                                                                                                                (Must be completed for Projects
Project                Description                                 Funding tive Sponsor         Manager       Status   Project Start   Project Finish   % Complete                                               Status                                                       Yellow or Red status)
Active Directory Kit   Develop a “Self Serve” kit for clients to      --     --   Barry M       Caren K       Yellow    5-Dec-05           TBD             5%      3/3: Meet with WST and Sponsor this week to determine project duration.                     2/24:    3/3: No change
                       migrate into Active Directory.                            for John F                                                                        No update on this project until the week of 3/6                                                      2/24: No change.




EMC Storage            Resolve all outstanding storage                --     --    John F        Steve L      Yellow       tbd              NA              NA      3/2: Disks are here, process for new BCV can start                                                  3/2: Resource constraints need
                       issues, including Forsythe SAN                                                                                                               3/2: Clarion Standby's are in progress, FW design complete                                          to be addressed, other tech
                       Replacement (Migration of Forsythe                                                                                                           3/2: 600B to D migration underway - 1 TSM instance complete                                         issues are solved or underway
                       SAN to Cisco equipment dual-                                                                                                                 2/24: New disks to arrive next week                                                                 2/24: Need firewall design for
                       attaching each host to eliminate single                                                                                                      2/17 & 2/10: New DMX mitigation using the Clarion is underway. Further refining of the "no          clarion
                       point of failure) and EMC Storage                                                                                                            dynamic spare" and new disk for new BCV options. 600B to D migration progress slowed due to         2/24: Need 5 day blackout
                       Strategy.                                                                                                                                    DMX effort and operational priorities                                                               period from AS


Loadbalancing          Global Load Balancing Service for              --     22    John F      Bernadette D   Yellow    1-Jan-06        31-Aug-06          10%      2/24: No significant progress.                                                                      2/10: Has been red on client
Service                Clients:                                                                                                                                     2/17: New direction to pilot F5 w/GSB and focus on long-term architecture and LB opportunities.     report for a number of weeks. It
                       Initiate load balancing service to                                                                                                           Requirements meeting w/GSB scheduled for 3/2. Project charter in re-write, will be finalized        should remain in Yellow until
                       enhance system performance and                                                                                                               after John F.'s vacation 3/1.                                                                       project charter is approved and
                       availability                                                                                                                                                                                                                                     timeline is set.




RWC - ECH &            Prepare Data Center in Redwood City            --     41     Jay K         Bob M       Yellow     1-Oct-05       30-Jun-07          10%      3/7: No Change in status from 2/21 update. 510 Broadway data center design is still in              3/7: outcome and timing of 510
Service                (RWC) for Hospital                                                                                                                           progress. This will be the last phase of the project design pending outcome of ITS/BA               Broadway data center pending
Provisioning for                                                                                                                                                    consultant review. The review is scheduled to be completed by June 2006. If an ECH is               ITS/BA consultant review.
Hospital                                                                                                                                                            required to support the Hospital prior to 510 being ready, an ECH will be installed within one of
                                                                                                                                                                    their buildings as an interim solution. This delay will effect the original complete date of Nov,
                                                                                                                                                                    2006.
TimeTracking           a. Standardize, reconfigure                    --     38    Sam S       Bernadette D   Yellow    1-Jan-06         1-Jun-06          10%      3/3: Project on hold while Sam, Bill & Jan review other application offers. Sponsor meeting         2/24: Resource still the key
Improvement            Timetracking application                                                                                                                     scheduled next week.                                                                                issue.
                       b. Ensure current categories and                                                                                                             2/24: Conf call w/Journyx to understand existing hierarchy and discuss upgrade                      2/17: Appl Admin/Analyst
                       configurations are useful to staff as                                                                                                        process/options.                                                                                    position is critical for this project
                       well as rate setting                                                                                                                                                                                                                             and has not been staffed yet. If
                       c. Educate / motivate managers to use                                                                                                                                                                                                            staffing doesn't occur soon, this
                       application for staff resource                                                                                                                                                                                                                   project will be delayed or has to
                       management and future planning                                                                                                                                                                                                                   be limited in scope.

Unified messaging      Pilot to determine replacement of              --     --     Jay K       Michelle C    Yellow        ??              ??              0%      3/3 - No additional updates, waiting for project approval.                                          2/17: Need to know if this
Pilot                  existing OCTEL Voice Mail                                                                                                                    2/17: Conference call with AVST vendor and Sean Riordan to determine the scope of his               project is a priority and if there is
                       messaging system.                                                                                                                            involvement - min of 240 hours over the 6 month pilot, not including team meetings time             commitment to moving this
                                                                                                                                                                    2/10: Project will be delayed until April-May timeframe as necessary resources are not available    forward. This decision needs to
                                                                                                                                                                    at this time.                                                                                       be made soon as vendors need
                                                                                                                                                                                                                                                                        to be notified of the decision.

                                                                                                                                                                                                                                                                        WSG staff are critical for this
                                                                                                                                                                                                                                                                        project . Due to their staff
                                                                                                                                                                                                                                                                        shortage and existing
Order Management Business Process Redesign for Order                  --      8     Jay K        BPR =         Red      1-Sep-05        31-Aug-06          10%      3/3:Team Documentation submitted to Sponsor for review. Steering Committee Meeting to be            2/10: Change in scope to include
BPR              Management process. Originally                                                 Christine                                                           scheduled to review results, decisions and implementation plan. 2/24: Daily team meetings           Data center.
                 scoped to Telecom orders; 12/2005                                               Yelda                                                              concluded on 2/22. Completed team objectives. Team is completing documentation and will
                 added Data Center billing as well.                                                                                                                 make final edits by 2/28 and submit to Sponsor 3/1.
                 Determine and implement revised
                 processes and systems to support
                 them.

Equipment            Design and implement equipment                   --      9    John F       BPR = Liz     Moved     1-Sep-05        31-Aug-06          10%      2/24: Project incorporated into Order Management BPR
Provisioning BPR     provisioning process for servers. This                                    Goesseringe
                     includes ordering, database, system                                            r
                     administration, firewall.
Continuing Studies - Transition Continuing Studies Web &           Client    --     Barry M      Caren K      Green     1-Feb-06         10-Apr-06         20%      3/3: Had meeting with CSP, security, and PCI. Will be working on the remediation plan for PCI.
eCommerce            database server from a 1 tier windows                        for John F                                                                        2/24: Kickoff meeting last week. Working with the client regarding the PCI compliance meeting.
Implementation       server environment to a more secure,                                                                                                           WST is working on a test upgrade.
                     PCI compliant 2 tier windows server
                     environment with firewalls.




        11/30/2010                                                                                                                                                                                                                                                                     page 1 of 6
IT Services Weekly Project Status Report
6-Mar-06



                                                                                                                                                                                                                                                                                  Hot Issues
                                                                         Initia Project       Project                                                                                                                                                                   (Must be completed for Projects
Project              Description                                 Funding tive Sponsor        Manager      Status   Project Start   Project Finish   % Complete                                            Status                                                            Yellow or Red status)
Desktop Backup       Transition PC desktop to third-party           --     23   Jan C        Caren K      Green      1-Apr-05       31-Dec-06          50%     3/3: Setting up communication plan meetings with various groups, completed e-mail write up,
                     vendor for desktop backup. Update                                                                                                         and Connected home page message. Will be reviewing all documentation with the group.
                     Mac desktop backup on TSM.                                                                                                                2/24: Met with Sponsor. Making progress on the communication plan requirements


Active Directory     Assist specific clients with migration to     --      --    Barry M      Caren K     Green     1-Sep-05        31-Aug-06       LAW: 5%  3/3: OOD: Worked with OOD to troubleshoot logon scripts. Things are progressing as
Migrations           Active Directory.                                          for John F                                                           OOD 30% planned.                                                                        2/24: OOD
                                                                                                                                                             Met with OOD last week. The have successfully migrated computers into the test environment.
                                                                                                                                                             Project is on schedule.                                                   3/3: No update
                                                                                                                                                             on the Law project.                                             2/24: LAW: No updates
                                                                                                                                                             since last week. Tim Poston is working on the final contact.

AMCOM Upgrade        Upgrade AMCOM paging system.                  18      --     Jay K      Michelle C   Green     15-Nov-05       31-Mar-06          75%       3/3: AMCOM has resolved the Oracle-AMCOm app problem - we are on target for the cutover
                                                                                                                                                                 on the 21st.

                                                                                                                                                                 2/24:
                                                                                                                                                                 1) Communications completed, including text for AMCOM website and login screen.
                                                                                                                                                                 2) Completed planning of staged upgrade of operator PCs, including OS and app for March 7th.
                                                                                                                                                                 3) Training venue scheduled for OnCall web admins, March 1-3.
                                                                                                                                                                 4) AMCOM coming this week to train all 28 operators, SMC and paging staff
                                                                                                                                                                 5) Process defined for distributing the passwords.
                                                                                                                                                                 2


Audit Response -     Remediate security audit findings             --     32      Bill C      Caren K     Green     5-Aug-06        31-Aug-06          80%       3/3: Final revision of the "Category A Data" Infrastructure and HATS Care Data item will be
Internal                                                                                                                                                         completed within the next few week. Also, met with Sponsor this week. Will work on a plan to
                                                                                                                                                                 track and corrdinate if nessasary the tasks that are created due to the audit responses.
                                                                                                                                                                                                                                                2/24: Steven
                                                                                                                                                                 submitted his document. Waiting to hear back from Eric N if this will satisfy the requirement to
                                                                                                                                                                 close this item. Momentum has slowed down for the "Category A" Data response that ITS will
                                                                                                                                                                 submit a piece for. Bruce has a draft write up. The question if the ED's want us to submit the
                                                                                                                                                                 draft as the final response.

Audit Response -     Coordinate IT Services responses for          --     33      Jay K       Caren K     Green     5-Aug-06        31-Aug-06        100% for    3/3: Met with the Sponsor. Will be working on the coordination efforts of the work that needs to
External (PWC)       recommendations from PWC.                                                                                                       the 2005    be completed from the 2005 and prior responses.                                            2/24: No
                                                                                                                                                    Responses    updates.
Change               Implement additional requirements for         --     20      Jay K        Bill H     Green     1-Jan-06        31-Mar-06           90%      3/6: The first two phases of the production rollout occurred on 2/28 and 3/3. Approximately 40
Management           Change Management tools and                                                                                                                 Change Requests have been submitted without major incident. The next phase occurs on 3/8,
Phase 2              processes. Complete phased rollout                                                                                                          with a final phase on 3/20. All but one of the scheduled training sessions have been held, with
                     to IT Services and Administrative                                                                                                           roughly half of the registered Change Management users having attended. Additional sessions
                     Systems.                                                                                                                                    will be scheduled based on demand from the user community.
                                                                                                                                                                 2/24: User training is underway. The rollout to full production is still scheduled to start on 2/28.


Desktop             Support the application and                   SGG     14     Barry M      Steve L     Green     1-Jan-06         1-Mar-06          80%       3/2: Working on production handoff, and governance process                                             3/2: Need to setup production
Configuration Mgmt- maintenance of Category A (HIPAA)                           for John F                                                                       2/24: Web site is almost complete, new product owner is engaged in onoging processes                   governance processes
Phase 2             required security settings, as well as
                    allow local customization of the DCM
                    profiles, allowing local IT groups to
                    tailor the profile to their needs.

Disaster          - Review, prioritize, and fulfill highest       Jung    43     John F       Steve L     Green     1-Sep-05         1-Sep-06          40%       3/2: JohnF and SteveJ connected, reviewed near term work with IBM                                      3/2: IT Finance needs to confirm
Recovery/Business priority EOC data needs                                                                                                                        2/24: IBM proposal in house, review is underway                                                        project scope and funding with
Continuity        - Determine disaster recovery                                                                                                                  2/24: Final interview results still pending                                                            Noel and SteveJ
                  strategies for information deemed to                                                                                                                                                                                                                  3/2: EOC members need to
                  be of highest priority to the                                                                                                                                                                                                                         respond
                  University’s emergency response                                                                                                                                                                                                                       3/2: IBM meeting on 3/6 to
                  - Implement one or more business                                                                                                                                                                                                                      review and revise scope
                  continuity services partnerships                                                                                                                                                                                                                      2/24: Confirm :smeaker net : for
                  - Continue detailed planning for                                                                                                                                                                                                                      April EOC exersize
                  disaster recovery and business
                  continuity




       11/30/2010                                                                                                                                                                                                                                                                    page 2 of 6
IT Services Weekly Project Status Report
6-Mar-06



                                                                                                                                                                                                                                                                            Hot Issues
                                                                         Initia Project     Project                                                                                                                                                               (Must be completed for Projects
Project               Description                                Funding tive Sponsor      Manager       Status   Project Start   Project Finish   % Complete                                                Status                                                    Yellow or Red status)
Firewall Operations   Project management of new clients to          --     --    Jay K     Michelle C    Green     30-Oct-05           NA             NA      3/3 - Coordinated meeting with Ty Parker (RDE); also after hours support for their db migration.
Management -          the firewall service, additions of zones                                                                                                2) prep for meeting with Larry Dillard re: Delphi firewall
Client & Other        to existing firewall projects and other                                                                                                 3) Coordinating with CR and Continuing Studies on meeting to discuss their fw implementation.
Projects              client firewall initiatives.                                                                                                            4) Created analysis of Delphi rules and fw topology, sent to Larry Dillard.
                                                                                                                                                              5) Assisted Change Management staff in understanding firewall processes, rules analysis and
                                                                                                                                                              resolving their inability to access a server.

                                                                                                                                                                2/24:
                                                                                                                                                                1) Coordinated meeting for Gerry Weitz, Bruce Seidel and Barry Magsany re: their need for
                                                                                                                                                                hosting and firewall.
                                                                                                                                                                2) Coordinated meeting with the TSM team, Steve Loving and the FWT for brain storming
                                                                                                                                                                about options for resolving the Clarion firewall issue.
                                                                                                                                                                2) Weekly meeting with Liz G on upcoming client projects.




Firewall Operations   A series of discrete projects focused        --      --    Jay K     Michelle C    Green     30-Oct-05           NA              NA       3/3
Management -          on the improvement of the operational                                                                                                     1) "Move and add a system docs" updated in conjunction with staff - additional re-writes
Operations            processes associated with the firewall                                                                                                    necessary.
Process               service. This includes rates                                                                                                              2) Met with Jane Marcus and Alvin Chew to review the SSP and PAC docs for N2Switch.
Development           development and associated tasks                                                                                                          3) Compiling docs for FW PAC review.
                      such as developing tools to automate
                      the metrics capturing capabilities for
                      billing purposes on the firewall,                                                                                                         2/24: Started prep for process review and editing of docs with staff.
                      upgrading existing client facing
                      documentation and processes, etc.

Firewall Operations Develop process for capturing,                 --      --    Jay K     Michelle C    Green     30-Oct-05           NA              NA       3/3: No new information
Management -        formalizing, and rolling out rates for                                                                                                      2/14: Met with Fred and Cholada to input PTAs for firewall billing
Rate Development firewall service.

HIPAA Secure          Pilot secure email service that allows     Provost   31   John F      Steve L      Green     1-Jan-05         1-Jun-06          70%       3/2: Still testing for both the move (need LB) and for Hi Vol bugs                                3/2: Still need stabilization,
Email                 Stanford users to send encrypted                                                                                                          2/24: Found performance issues, plan underway to move servers out of Forsythe                     moving ahead on funding need
                      email and mitigate HIPAA requirement                                                                                                      2/24 Found JAVA memory error on Hi Vol group                                                      for ongoing services
                      to facilitate encrypted email.                                                                                                                                                                                                              2/24: Need to move servers out
                                                                                                                                                                                                                                                                  of forsythe and stabilize service
                                                                                                                                                                                                                                                                  2/24: Need to solve the Hi Vol
                                                                                                                                                                                                                                                                  issues


Host Self             Implement automated method to               SGG      30   Jan C       Steve L      Green     1-Sep-05         1-Jun-06          42%       3/2: All on track, working on emerging issues of governance and tool maintenance                  3/2: May need to change scope
Registration          verify machines are not at risk prior to                                                                                                  2/24: Progress on wireless registration shown in todays meeting                                   for WSUS server to fix some
                      joining Stanford network. Ports                                                                                                           2/24: Identified barriers in LB area                                                              issues with patches and
                      existing implementation to java.                                                                                                          2/17: Met with LNA group for first marketing, all tasks on target                                 updates for new machines
                                                                                                                                                                                                                                                                  3/2: Need updated status on
                                                                                                                                                                                                                                                                  policy based routing (see 2/3
                                                                                                                                                                                                                                                                  issue)
                                                                                                                                                                                                                                                                  2/17: HCT may be impacted by
                                                                                                                                                                                                                                                                  new "checker" sw support from
                                                                                                                                                                                                                                                                  MS
                                                                                                                                                                                                                                                                  2/3: Significant dependencies
                                                                                                                                                                                                                                                                  on NetDB and Network work still
                                                                                                                                                                                                                                                                  exist

Inventory Database Inventory Database                              --      --    Jay K    Bernadette D   Green     1-Sep-05        31-Aug-06          55%       2/24: Updated documentation with latest changes. In process of validating the latest data load.
                                                                                                                                                                Documented and reviewed specifications for ongoign data loads with Unix team. Interface code
                                                                                                                                                                being developed. Rollout meeting with key stakeholders scheduled for 3/3.
                                                                                                                                                                2/17: Additional changes have been made. Are in process of loading latest data using
                                                                                                                                                                spreadsheets submitted to Finance for Pinnacle Feb run.

Kerberos Migration    Retire Stanford's Kerberos 4 cell and       SGG      34   John F       Jon P       Green     1-Jan-06         1-Apr-08          50%       3/4: UI finalized for weblogin. Webauth documentation updated. Tony published a
(4 --> 5)             move all centrally-provided services                                                                                                      requirements doc for ESD, to be reviewed by community. Confirmed ESD will not support
                      that rely on Kerberos 4 for                                                                                                               Win2K—they can continue ot use PCLeland. We expect to hear back from Sine Nomine this
                      authentication to use Kerberos 5.                                                                                                         week with a proposal. 2/27: Released Webauth 3.4.2 (bugfix). UI reviews almost done. Still
                                                                                                                                                                on target for 3/15 PC OpenAFS and 4/1 Weblogin release. Mac OpenAFS pending upstream
                                                                                                                                                                release of 1.4.1




       11/30/2010                                                                                                                                                                                                                                                              page 3 of 6
IT Services Weekly Project Status Report
6-Mar-06



                                                                                                                                                                                                                                                                                    Hot Issues
                                                                        Initia Project         Project                                                                                                                                                                    (Must be completed for Projects
Project               Description                               Funding tive Sponsor          Manager       Status     Project Start   Project Finish   % Complete                                                   Status                                                    Yellow or Red status)
Outsource             Goal: work with clients to identify          --     --    Jay K         Michelle C    Green        1-Oct-05       31-May-06          80%     3/3:
Highspeed Printing    suitable alternatives to Forsythe                                                                                                            1)Last client transition meeting with Eric Armstrong.
                      printing, assist clients with the                                                                                                            2) Remaining data provided to Procurement for RFP
                      transition to the new method, and                                                                                                            3) Meeting planned with the Controllers office for 3/23
                      shutdown the forsythe print service by                                                                                                       2/21:
                      6/1/06.                                                                                                                                      1) Erik C and I met with Sue Webert and Juliet Harris from Procurement re: the RFP for
                                                                                                                                                                   printing. I will provide the data for this by 2/28.
                                                                                                                                                                   2) Erik C and I met with Jayshri and Raj from the CO re: the PGP encryption of their data. They
                                                                                                                                                                   have a new need to have their data archived to CD and encrypted. We are following up
                                                                                                                                                                   internally to see if this is a service for fee that we can provide.
                                                                                                                                                                   3) Have scheduled meetings with all clients to discuss their printing transition.

PMO BPR               Develop a set of processes and tools         --      7      Jan C        Joyce D      Green        5-Oct-05         6-Aug-06         60%       3/1: Finance pre-meeting resolved out standing issues. Revised Charter created & distributed.
                      to more effectively manage projects in                                                                                                          Embedded PM Tracking sheet in progress. Resource tracking sheet created and reviewed.
                      IT Services.                                                                                                                                   Website Draft released.
                                                                                                                                                                     2/27: Weekly status report well received. Focus on implementing project selection & staffing
                                                                                                                                                                     process. Updating templates. Will verify finance tracking with next set of financial statements.




RWC - Data Center Develop integrated strategy for                  --      40    John F        Jon P        Green        1-Sep-05         1-Jun-07         10%       3/4: Meeting with Bill later this month. No other update. 2/27: Canceled remaining Service           2/17: Need to clarify PM role
Strategy          datacenter.                                                                                                                                        Move meetings until post discovery phase. Setting up periodic meetings w/Bill and John to            going forward
                                                                                                                                                                     track discovery phase.

Shibboleth         The project will deploy a Shibboleth          SGG       35    John F        Jon P        Green        1-Jan-06        31-Aug-06          5%       3/3: Bruce is running InCommon application by legal (he wasn't sure about an indemnity clause
                   infrastructure that will support the                                                                                                              in the contract). Packaging work started (Quanah). Agreed to try to rollout shib on
                   authentication of Stanford users to                                                                                                               www.stanford.edu as a test area. Looking to create Stanford federation as part of “friends and
                   access authorized online web                                                                                                                      family” SUNetID project (proposal) for next year. 2/27: mod_shib packaging work delayed
                   resources at other participating                                                                                                                  slightly. Scotty to present chalk talk this Friday. InCommon and Admin Guide work ongoing.
                   institutions. The project will also
                   provide documentation for Stanford
                   service providers who want to make
                   their web applications available to
                   users across federations.
Sybase Elimination Complete the port of NetDB and                SGG       15     Jay K        Jon P        Green        1-Jan-05         1-Jun-06         70%       3/3: Developers took their last piece of training. Gaby started QA with Sunia. Alpha6 (QA
for NetDB          related technologies from Sybase to                                                                                                               release) code complete. from the front-end. 2/27: Quality assurance position filled, starts 3/1 or
                   Oracle, and refactor the database                                                                                                                 3/15 (still clarifying). QA cycle in progress.
                   interface layer of the NetDB system.
Change             Pilot new Change Management                     --      19     Jay K         Bill H     Complete      1-Mar-05        31-Dec-05         100%      Done. Pilot live 12/14/05.
Management         process and system to improve IT
Phase 1            Services’ ability to track, approve, and
                   implement changes to technologies,
                   architecture and service levels. Pilot
Desktop            CRC desktop configuration for all of          SGG       14     Jan C        Steve L     Complete      1-Mar-05         1-Dec-05         100%      Done.
Configuration Mgmt CRC clients.


Status is defined as the Percentage Complete, and then Green, Yellow or Red:
     Complete                                            The project has been closed and is 100% complete.
        Hold                                             The project is postponed or cancelled. No work is being done at the time.
       Green                                             The project is on track to be completed as scheduled and within scope and budget. Confidence level is 90% or greater.
      Yellow                                             The project is at risk. It should still hit its completion targets, but there are known problems that threaten to take it off course. Confidence level is 70-90%.
         Red                                             The project is off track. There have already been problems impacting completion or they can been seen on the horizon. Confidence level is below 70%.

Descriptions of the individual projects can be found at:       www.stanford.edu/dept/its/organization/projectmanagement

To receive this report weekly from the IT Services Project Management Office, subscribe to:                its-projectstatus@lists.stanford.edu




        11/30/2010                                                                                                                                                                                                                                                                     page 4 of 6
ResComp                       Jan C   Michelle C   TBD     TBD      0%      Green     1/23 :Not
 Security                                                                             yet a
Interventi                                                                            project,
on and                                                                                probably
Remediati                                                                             will result
on system                                                                             in an
                                                                                      initial
                                                                                      meeting,
                                                                                      and some
                                                                                      consolidat
                                                                                      ion of
                                                                                      existing
                                                                                      web
                                                                                      based




RWC -        Develop     --    45       Jay K      Jon P   Hold   1-Sep-05 1-Jun-07      5%
Staff        a
Relocatio    tentative
n            plan for
             the move
             of IT
             Services
             employee
             s to
             Redwood
             City.
1/30: On
hold until
staff
move
strategy
defined.
Staff
moves
related to
Data
Center
move
incorporat
ed in
RWC-
Services
Move
project.

				
DOCUMENT INFO
Description: 3.15 Project Management What Is the Project Duration in Weeks document sample